OpenStudy (anonymous):

Write the equation of the line that passes through (1, 5) and (–2, 14) in slope-intercept form. y = 3x + 2 y = 3x + 8 y = –3x – 2 y = –3x + 8

OpenStudy (texaschic101):

first we have to find the slope by using the slope formula slope(m) = (y2 - y1) / (x2 - x1) (1,5)....x1 = 1 and y1 = 5 (-2,14) ...x2 = -2 and y2 = 14 now we sub slope(m) = (14 - 5) / (-2 - 1) slope(m) = 9/-3 slope(m) = -3 now we use y = mx + b we now know that the slope(m) = -3 you can use either of your points...I will use (1,5)...x = 1 and y = 5 now all we have to solve for is b, the y intercept so lets solve for b y = mx + b 5 = -3(1) + b 5 = -3 + b 5 + 3 = b 8 = b so our equation is : y = -3x + 8

OpenStudy (anonymous):

thank you soo much if u have time can u help me with 2 more problems bcz im lost

OpenStudy (texaschic101):

I can try :)

OpenStudy (anonymous):

thank you so much. your a life saver

OpenStudy (anonymous):

Anne has a total of 3,500 golf balls at home. She loses some golf balls when she goes out on the driving range daily. The number of golf balls, h(x), can be represented by the function h(x) = –150x + 3,500, where x is the number of days she plays golf. Suppose 3,500 changed to 4,000. What does this mean in the context of the problem? Changing 3,500 to 4,000 means the initial number of golf balls decreased. Changing 3,500 to 4,000 means the initial number of golf balls increased. Changing 3,500 to 4,000 means the daily number of golf balls lost decreased. Changing 3,500 to 4,000 means the daily number of golf balls lost increased.

OpenStudy (anonymous):

Write the equation of the line that passes through (–1, 5) and has a slope of 3 in point-slope form. x + 1 = 3(y – 5) x – 5 = 3(y + 1) y – 5 = 3(x + 1) y + 1 = 3(x – 5)

OpenStudy (texaschic101):

Initial number of golf balls increased. you see, she started with 3500...and it went up to 4000 ===================== (-1,5)....slope(m) = 3 now we use point-slope formula y - y1 = m(x - x1) (-1,5).....x1 = -1 and y1 = 5 m = 3 y - 5 = 3(x - (-1) y - 5 = 3(x + 1)

OpenStudy (anonymous):

so for the first question it would be B

OpenStudy (texaschic101):

yes

OpenStudy (anonymous):

THANK YOU SO MUCH. lol i would give u a big hug right now

OpenStudy (texaschic101):

lol.....sweet...glad to help :)

OpenStudy (anonymous):

lol is there anyway i can contact you on here because i suck at algebra :(

OpenStudy (texaschic101):

you can try to tag me...but keep in mind, there are several people on here that are better then I am at this stuff...you just happened to ask about something I knew about.

OpenStudy (anonymous):

oo ok. lol but thank you so much thx to you i got 100% on my virtual school assignment

OpenStudy (texaschic101):

sweet....doing the happy dance for you now :)